I need help with this equation : 8+-2(x+5)-(x+4)=3x+(-8)+1^77
Answer by Alan3354(69443)

8+-2(x+5)-(x+4)=3x+(-8)+1^77
8+-2(x+5)-(x+4)=3x+(-8)+1
8+-2(x+5)-(x+4)=3x - 7
8 -2(x+5)-x-4 = 3x - 7
8 - 2x - 10 - x - 4 = 3x - 7
8 - 3x - 14 = 3x - 7
-6 - 3x = 3x - 7
-6 = 6x - 7
1 = 6x
x = 1/6